OpenStudy (anonymous):

simplify the following quotient of complex numbers into the form a+bi. -4-3i/2-i

OpenStudy (anonymous):

a=-1 b=-2

OpenStudy (phi):

Multiply top and bottom by the complex conjugate of (2-i) (which is 2+i ) I'll do the bottom: (2-i)(2+i)= 2*2- 2*i + 2*i -i*i= 4 +1 = 5